The probability that a city bus is ready for service when needed is 83%. The probability that a city bus is ready for service and has a working radio is 68%. Find the probability that a bus chosen at random has a working radio given that it is ready for service. Round to the nearest tenth of a percent.
Answer by stanbon(57377)

The probability that a city bus is ready for service when needed is 83%.
---
P(ready) = 0.83
---------------------------
The probability that a city bus is ready for service and has a working radio is 68%.
---
P(ready and radio) = 0.68
----------------------------
Find the probability that a bus chosen at random has a working radio given that it is ready for service. Round to the nearest tenth of a percent.
---
P(radio|ready) = P(radio and ready)/P(ready) = 0.68/0.83 = 0.82
============================
Cheers,
Stan H.